Let m= Maricar's age now. Let b= Bong's age now. Express the following in terms of m and b.
a. Bong's age in 7 years
b. Bong's age 4 years ago
c. Maricar's age in 7 years
d. Maricar's age 4 years ago
e. The sum of their ages in 7 years
f. The sum of their ages 4 years ago

Answer by Maths68(1474)

a. Bong's age in 7 years=b+7
b. Bong's age 4 years ago=b-4
c. Maricar's age in 7 years=m+7
d. Maricar's age 4 years ago=m-4
e. The sum of their ages in 7 years
=(b+7)+(m+7)
=b+m+14
f. The sum of their ages 4 years ago
=(b-4)+(m-4)
=b+m-8